# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v7

@@ IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
TABLE III: Well-designed methods in the ICL paradigm for LLM-based text-to-SQL are ordered by release time. The methods are grouped into four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement. Methods that fall into multiple categories will be introduced accordingly. The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [yu2018Spider], BIRD [li2023BIRD], Spider-DK [gan2021Spider-DK] Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], and Spider 2.0 [lei2025spider2]
. We use the reference ID in the table due to column space limitations. *There are multiple applied LLMs in the corresponding method; the model with representative performance is presented. CoT method is introduced in multiple venues: NeurIPS’23 [li2023BIRD], EMNLP’23 [tai2023exploring], VLDB’24 [gao2023dailsql], arXiv’24 [zhang2024sgusql]. Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue MBR-Exec [shi2022natural] Codex [yu2018Spider] EX Apr-2022 EMNLP’22 Coder-Reviewer [zhang2022coder] Codex [yu2018Spider] EX Nov-2022 ICML’23 LEVER [ni2023lever] Codex [yu2018Spider] EX Feb-2023 ICML’23 Self-Debugging [chen2024selfdebugging] StarCoder* [yu2018Spider] EX Apr-2023 ICLR’24 DESEM+P [guo2023desem] ChatGPT [yu2018Spider, gan2021Spider-SYN] EX Apr-2023 PRICAI’23 DIN-SQL [pourreza2023dinsql] GPT-4* [li2023BIRD, yu2018Spider, lei2025spider2] EX, EM, VES Apr-2023 NeurIPS’23 CoT [tai2023exploring, li2023BIRD, gao2023dailsql, zhang2024sgusql] GPT-4 [yu2018Spider, deng2021Spider-Realistic, li2023BIRD] EX, VES May-2023 Multiple Venues StructGPT [jiang2023structgpt] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 SD+SA+Voting [nan2023enhancing] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Findings QDecomp [tai2023exploring] Codex [yu2018Spider, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Least-to-Most [tai2023exploring] Codex [yu2018Spider] EX May-2023 EMNLP’23 SQL-PaLM [sun2023sqlpalm] PaLM-2 [yu2018Spider] EX May-2023 TMLR’24 RAG+SP&DRC [guo2023ragsql] ChatGPT [yu2018Spider] EX Jul-2023 ICONIP’23 C3 [dong2023c3] ChatGPT [yu2018Spider] EX Jul-2023 arXiv’23 DAIL-SQL [gao2023dailsql] GPT-4* [yu2018Spider, li2023BIRD, deng2021Spider-Realistic, lei2025spider2] EX, EM, VES Aug-2023 VLDB’24 ODIS [chang2023selective] Codex* [yu2018Spider] EX Oct-2023 EMNLP’23 Findings ACT-SQL [zhang2023actsql] GPT-4* [yu2018Spider, gan2021Spider-SYN] EX, EM Oct-2023 EMNLP’23 Findings MAC-SQL [wang2024macsql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Dec-2023 COLING’25 DEA-SQL [xie2024deasql] GPT-4 [yu2018Spider] EX Feb-2024 ACL’24 Findings FUSED [wang2024fused] ChatGPT* [yu2018Spider] EX Feb-2024 EMNLP’24 Findings DELLM [hong2024knowledgetosql] GPT-4* [yu2018Spider, li2023BIRD] EX, VES Feb-2024 ACL’24 Findings SGU-SQL [zhang2024sgusql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM Feb-2024 ICML’25 PoT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24 SQL-CRAFT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24 FUXI [gu2024fuxi] GPT-4* [li2023BIRD] EX Feb-2024 EMNLP’24 MetaSQL [fan2024metasql] GPT-4* [yu2018Spider] EX, EM Feb-2024 ICDE’24 $\textit{{R}}^{3}$ [xia2024r3] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 ACL’25 Workshop PET-SQL [li2024petsql] GPT-4 [yu2018Spider] EX Mar-2024 arXiv’24 PURPLE [ren2024purple] GPT-4* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX, EM Mar-2024 ICDE’24 Dubo-SQL [thorpe2024dubo] ChatGPT* [li2023BIRD] EX Apr-2024 arXiv’24 TA-SQL [qu2024generation] GPT-4* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] EX Apr-2024 ACL’24 Findings MCS-SQL [lee2025mcs] GPT-4 [yu2018Spider, li2023BIRD] EX, VES May-2024 COLING’25 CHESS [talaei2024chess] GPT-4o* [yu2018Spider, li2023BIRD, lei2025spider2] EX, VES May-2024 arXiv’24 SuperSQL [li2024dawn] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Jun-2024 VLDB’24 Distillery [maamari2024the] GPT-4o-Mini* [li2023BIRD] EX Aug-2024 NeurIP’24 Workshop E-SQL [caferouglu2024esql] GPT-4o* [yu2018Spider, li2023BIRD] EX, VES Sep-2024 arXiv’24 CHASE-SQL [pourreza2025chasesql] Gemini-1.5* [yu2018Spider, li2023BIRD] EX, VES Oct-2024 ICLR’25 RSL-SQL [cao2024rsl] DeepSeek-V3* [yu2018Spider, li2023BIRD] EX, VES Nov-2024 arXiv’24 Spider-Agent [lei2025spider2] o1-preview* [lei2025spider2] EX Nov-2024 ICLR’25 ICRL [toteja2025context] ChatGPT [yu2018Spider] EX Jan-2025 COLING’25 Gen-SQL [shi2025gen] LLaMA-3* [yu2018Spider, li2023BIRD] EX, EM, VES Jan-2025 COLING’25 SAFE-SQL [lee2025safe] GPT-4o-Mini* [yu2018Spider] EX, EM Feb-2025 arXiv’25 ReFoRCE [deng2025reforce] o1-preview* [lei2025spider2] EX Feb-2025 ICLR’25 Workshop LinkAlign [wang2025linkalign] DeepSeek-R1* [li2023BIRD, lei2025spider2] EX, EM Mar-2025 arXiv’25

### v8

@@ IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
TABLE III: Well-designed methods in the ICL paradigm for LLM-based text-to-SQL are ordered by release time. The methods are grouped into four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement. Methods that fall into multiple categories will be introduced accordingly. The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [yu2018Spider], BIRD [li2023BIRD], Spider-DK [gan2021Spider-DK] Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], and Spider 2.0 [lei2025spider2]. We use the reference ID in the table due to column space limitations. *There are multiple applied LLMs in the corresponding method; the model with representative performance is presented. CoT method is introduced in multiple venues: NeurIPS’23 [li2023BIRD], EMNLP’23 [tai2023exploring], VLDB’24 [gao2023dailsql], arXiv’24 [zhang2024sgusql].
Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
MBR-Exec [shi2022natural] Codex [yu2018Spider] EX Apr-2022 EMNLP’22
Coder-Reviewer [zhang2022coder] Codex [yu2018Spider] EX Nov-2022 ICML’23
LEVER [ni2023lever] Codex [yu2018Spider] EX Feb-2023 ICML’23
Self-Debugging [chen2024selfdebugging] StarCoder* [yu2018Spider] EX Apr-2023 ICLR’24
DESEM+P [guo2023desem] ChatGPT [yu2018Spider, gan2021Spider-SYN] EX Apr-2023 PRICAI’23
DIN-SQL [pourreza2023dinsql] GPT-4* [li2023BIRD, yu2018Spider, lei2025spider2] EX, EM, VES Apr-2023 NeurIPS’23
CoT [tai2023exploring, li2023BIRD, gao2023dailsql, zhang2024sgusql] GPT-4 [yu2018Spider, deng2021Spider-Realistic, li2023BIRD] EX, VES May-2023 Multiple Venues
StructGPT [jiang2023structgpt] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23
SD+SA+Voting [nan2023enhancing] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Findings
QDecomp [tai2023exploring] Codex [yu2018Spider, deng2021Spider-Realistic] EX May-2023 EMNLP’23
Least-to-Most [tai2023exploring] Codex [yu2018Spider] EX May-2023 EMNLP’23
SQL-PaLM [sun2023sqlpalm] PaLM-2 [yu2018Spider] EX May-2023 TMLR’24
RAG+SP&DRC [guo2023ragsql] ChatGPT [yu2018Spider] EX Jul-2023 ICONIP’23
C3 [dong2023c3] ChatGPT [yu2018Spider] EX Jul-2023 arXiv’23
DAIL-SQL [gao2023dailsql] GPT-4* [yu2018Spider, li2023BIRD, deng2021Spider-Realistic, lei2025spider2] EX, EM, VES Aug-2023 VLDB’24
ODIS [chang2023selective] Codex* [yu2018Spider] EX Oct-2023 EMNLP’23 Findings
ACT-SQL [zhang2023actsql] GPT-4* [yu2018Spider, gan2021Spider-SYN] EX, EM Oct-2023 EMNLP’23 Findings
MAC-SQL [wang2024macsql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Dec-2023 COLING’25
DEA-SQL [xie2024deasql] GPT-4 [yu2018Spider] EX Feb-2024 ACL’24 Findings
FUSED [wang2024fused] ChatGPT* [yu2018Spider] EX Feb-2024 EMNLP’24 Findings
DELLM [hong2024knowledgetosql] GPT-4* [yu2018Spider, li2023BIRD] EX, VES Feb-2024 ACL’24 Findings
SGU-SQL [zhang2024sgusql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM Feb-2024 ICML’25
PoT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24
SQL-CRAFT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24
FUXI [gu2024fuxi] GPT-4* [li2023BIRD] EX Feb-2024 EMNLP’24
MetaSQL [fan2024metasql] GPT-4* [yu2018Spider] EX, EM Feb-2024 ICDE’24
$\textit{{R}}^{3}$ [xia2024r3] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 ACL’25 Workshop
PET-SQL [li2024petsql] GPT-4 [yu2018Spider] EX Mar-2024 arXiv’24
PURPLE [ren2024purple] GPT-4* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX, EM Mar-2024 ICDE’24
Dubo-SQL [thorpe2024dubo] ChatGPT* [li2023BIRD] EX Apr-2024 arXiv’24
TA-SQL [qu2024generation] GPT-4* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] EX Apr-2024 ACL’24 Findings
MCS-SQL [lee2025mcs] GPT-4 [yu2018Spider, li2023BIRD] EX, VES May-2024 COLING’25
CHESS [talaei2024chess] GPT-4o* [yu2018Spider, li2023BIRD, lei2025spider2] EX, VES May-2024 arXiv’24
SuperSQL [li2024dawn] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Jun-2024 VLDB’24
Distillery [maamari2024the] GPT-4o-Mini* [li2023BIRD] EX Aug-2024 NeurIP’24 Workshop
E-SQL [caferouglu2024esql] GPT-4o* [yu2018Spider, li2023BIRD] EX, VES Sep-2024 arXiv’24
CHASE-SQL [pourreza2025chasesql] Gemini-1.5* [yu2018Spider, li2023BIRD] EX, VES Oct-2024 ICLR’25
RSL-SQL [cao2024rsl] DeepSeek-V3* [yu2018Spider, li2023BIRD] EX, VES Nov-2024 arXiv’24
Spider-Agent [lei2025spider2] o1-preview* [lei2025spider2] EX Nov-2024 ICLR’25
ICRL [toteja2025context] ChatGPT [yu2018Spider] EX Jan-2025 COLING’25
Gen-SQL [shi2025gen] LLaMA-3* [yu2018Spider, li2023BIRD] EX, EM, VES Jan-2025 COLING’25
SAFE-SQL [lee2025safe] GPT-4o-Mini* [yu2018Spider] EX, EM Feb-2025 arXiv’25
ReFoRCE [deng2025reforce] o1-preview* [lei2025spider2] EX Feb-2025 ICLR’25 Workshop
LinkAlign [wang2025linkalign] DeepSeek-R1* [li2023BIRD, lei2025spider2] EX, EM Mar-2025 arXiv’25